FEL Upgrade Project Weekly Brief
December 4-8, 2000


We made further progress with harmonic conversion of FEL light to green and deep UV light.  See the Operations section of this report for further details.


A meeting was held this week with the DOE Contracting Officer and the ONR Contract Monitor which lead to completion of a draft Memorandum of Agreement (MOA) and Statement of Work (SOW) for the FY 01 completion phase contract for the FEL upgrade.

WBS 3 (Beam Physics):

Accelerator Design Revision 1.1 activities continued.  Linac module energy gains have been identified as an issue;
excessive asymmetry in these gains leads to focusing/energy mismatch at various points in the acceleration cycle, with
resulting degradation of beam envelope performance, chromatic aberration control, and beam transmission. Remedies are being investigated.

WBS 4 (Injector):

Vacuum baked the FET system with a titanium cathode.  The electrode gap was aligned and set at 1cm.  High voltage
conditioned the gap to 125 kV, 15pA baseline current.  Reset the gap to 6mm and will continue processing and

WBS 6 (RF):

Zone 4 - Completed preliminary checkout of the HPA and its interlocks.  Software has been written to test the zone.  As soon as the LCW is restored to the FEL, RF testing will start.

Zone 3 - No direct progress.  The documentation is being updated as a result of the problems found in zone 4, so zone 3
should come together quickly once parts are available.

WBS 8 (I&C):

This week we have <nearly> completed our move from the equipment room in the FEL to lab 5.  A hiring meeting is
scheduled for Monday for the coordinator position. This candidate, if they accept, will have been worth the wait! During
the process a candidate for the casual position was identified - this person is in transition from a NASA contract and is well suited to get the Laser Safety System upgrade on track.

The Optical Beam Position Monitor (OBPM) has beenn installed in the End-of-line (EOL) FEL beam dump this week. It is now obvious that a single attenuator for dynamic range is insufficient.  The latest opportunity came once we began
harmonic lasing at 1 micron - the 3 micro attenuator was opaque for that wavelength.  The optics group has purchased a 6 position color wheel to solve the color agility features.

Meetings were held to review the parts count for magnets and instrumentation for the upgrade and the UV add-on. The
viewers & BPMs were still consistant with this summer - the viewer contract is in progress and the BPM contract was
awarded this week. The FRPs were due 12/5/00 and turned out to be very competative - 6 vendors responded (from VA to NM) with quotes from $75k to $307k - most at low end with 2 shops each going an octave in $. Regarding trim magnet control the current approach is to combine all iron core magnets into the existing trim racks (utilizing the on/off hysteresis loop capabilities) and look for a lower cost bipolar supply for the air core magnets.  This could result in cost savings of $200k (90 air core with <$1k/ch vs. trim rack cost of ~$3k/ch). Commercial vendors are being consulted.

WBS 9 (Transport):

o We received budgetary estimates for our dipoles.
o  As a result of the magnetic modeling of the DW, we were able to assure ourselves that we could remove 20% of the
iron in the return legs and top and bottom poles while getting no degradation of the quality of the central field.  Some
additional cooling calculations convinced us that the copper cross section of this and the 180 degree dipole should be
increased.  These changes were reflected in the 3D drawing model and somewhat automatically flowed down to the details being prepared.  Additional GW dipole detailing continued.
o  We started the order for the copper for the GW Prototype.
QX (3.125" Quad)
o  Prototype Fabrication: Almost finished - the four quadrants were mechanically inspected and tolerances verified.
Interferences with the coil mounting were identified and corrected.  Core and coils were assembled and the bore was
verified to be in tolerance.  Plumbing and coil lead routing is almost complete.
o  Budgetary Estimates:  We are still awaiting three additional quotes from coil vendors.  These vendors were contacted
again, questions answered, and estimates should be received in the next few days.
o  Measurement Probe: Fabrication of the second analysis coil using 100-turn litz wire is progressing and should be finished next week.  Drawings for the probe body went through a design change to make assembly and inspection easier.
o  Comments on the modifications of the Differential Pumping stations to accommodate 3 inch tube were being assessed.
o  Dave Douglas notified us that the angles of the return lines in the Arcs are going to get smaller by about 5 degrees and that the pole face angles will change as he goes through his conversion of the lattice to an "Engineering" design.  We are ready to morph the values of the design spread sheet and the available magnet packages to the new values.
o  The Engineering Services Contract was signed by the two vendors and is in the DOE Office for approval.

WBS 10 (Wiggler):

Fabrication of the dispersion section is in progress.  A visit to the vendor is planned for the beginning of January.  The
drawings for the wiggler viewers were signed off.  The 3D design of the wiggler vacuum chamber support was finalized. The supplier of Group-3 Hall Probes (GMW) was contacted regarding information on planar hall effect specifications. They will fax the information.  D. Bullard attached the modified pole clamps to one of the wigglers.   Inspection of the gap showed a maximum spread of 0.0025" (spec 0.002".)   This shows that the design is acceptable.

WBS 11 (Optics):

We received bids for the mirror test stand vacuum vessel and the work has been awarded to a local vendor.  Work
proceeds at AES on the finite element analyses for the deformable mirror. The drawings for the internals of the mirror test stand are being reviewed.

We assisted in the installation of the holder for the quad pyro detector electronics for the O-BPM at the end-of-line
(EOL).  However, there are problems with the detector's dynamic range, and we and the I & C group are working on a
solution. Changes to some of the FEL diagnostics software were tested and debugged.


On Tuesday, beam was run for the Radiation Control group to verify radiation levels.  These measurements showed that the SRF cavities themselves were a source of neutron production. The RadCon group took data to determine whether radiation levels are sufficiently low to allow construction of the Helios addition without additional shielding.  Some FEL 3 micron beam was also given to the Nanotube research group for a run.

On Thursday the FEL was run for harmonic production studies.  We produced up to 26.6 W of average power at 524 nm using 100 W of IR light produced by third harmonic lasing. We then produced up to 2 W during a 230 microsecond
macropulse by doubling the 524 nm to 262 nm.  The overall efficiency was 5% from the IR to the UV.  The peak
efficiency seen for doubling was 37% using an LBO crystal.  The mode quality of the IR was rather poor so the overall
conversion efficiency can probably be increased. The micropulse length inferred from the spectrum was 0.5 psec for the IR.  It is noteworthy that we were able to produce the same level of 1 micron power as a month ago in about one hour of tuning.  More optimization is needed on the quadrupling however.  In addition, the fourth harmonic power is exquisitely sensitive to cavity length and cryomodule phasing which will complicate the delivery to users.

Privacy and Security Notice